OpenStudy (anonymous):

Algebra 2 Help, Please: Find the equation in standard form which passes through (–2, –7) and (1, –1). You must show all work.

OpenStudy (anonymous):

heres how i do these ones... i set it up in point slope first... so i find the slope to be rise/run which is (-1-(-7))/(1-(-2) which is 6/3 or 2....then i say y=2x+b..... to find b i just plug in 1 point... to get -1=2(1)+b..... so b=-3.... so then we have y=2x-3.... from there you should be able to put it into standard form!